# Yuehao Bai

### v7

@@ Abstract
This paper studies the use of finely stratified designs for the efficient estimation of a large class of treatment effect parameters that arise in the analysis of experiments. By a “finely stratified” design, we mean experiments in which units are divided into groups of a fixed size and a proportion within each group is assigned to a binary treatment uniformly at random. The class of parameters considered are those that can be expressed as the solution to a set of moment conditions constructed using a known function of the observed data. They include, among other things, average treatment effects, quantile treatment effects, and local average treatment effects as well as the counterparts to these quantities in experiments in which the unit is itself a cluster. In this setting, we establish three results. First, we show that under a finely stratified design, the naïve method of moments estimator achieves the same asymptotic variance as what could typically be attained under alternative treatment assignment mechanisms only through ex post covariate adjustment. Second, we argue that the naïve method of moments estimator under a finely stratified design is asymptotically efficient by deriving a lower bound on the asymptotic variance of regular estimators of the parameter of interest in the form of a convolution theorem. In this sense, finely stratified experiments are attractive because they lead to efficient estimators of treatment effect parameters “by design.” Finally, we strengthen this conclusion by establishing conditions under which a “fast-balancing” property of finely stratified designs is in fact necessary for the naïve method of moments estimator to attain the efficiency bound.
KEYWORDS: Convolution theorem, Efficiency, Experiment, Experimental design, Finely stratified experiment, Matched pairs, Randomized controlled trial
@@ 1 Introduction
This paper studies the use of finely stratified designs for the efficient estimation of a large class of treatment effect parameters that arise in the analysis of experiments. By a “finely stratified” design, we mean experiments in which units are divided into groups of a fixed size based on their covariate values and a proportion within each group is assigned to a binary treatment uniformly at random. The canonical example of such a design is a matched pairs design: here the fixed size of the groups equals two, and exactly one of the two units in each group is assigned to treatment at random, so that the marginal probability of treatment assignment is one half. More broadly, our use of the term “finely stratified” in this context contrasts with what we could call a “coarsely stratified” design, in which units are divided into a small set of large groups: see Example 4.2 below for a formal definition. The class of parameters considered are those that can be expressed as the solution to a set of moment conditions constructed using a known function of the observed data. This class of parameters includes many causal parameters of interest: average treatment effects (ATEs), quantile treatment effects, and local average treatment effects as well as the counterparts to these quantities in experiments in which the unit is itself a cluster.
@@
In the setting described above, we establish three results. First, we study the asymptotic properties of a naïve method of moments estimator under a finely stratified design. Here, by a naïve method of moments estimator, we mean an estimator constructed using a direct sample analog of the moment conditions. For example, in the case of the ATE, such an estimator is given by the Horvitz-Thompson estimator for the difference in means. We show that under a finely stratified design, the naïve method of moments estimator achieves the same asymptotic variance as what could typically be attained under alternative treatment assignment mechanisms only through ex post covariate adjustment using the same set of covariates. Such adjustment strategies frequently involve the nonparametric estimation of conditional expectations or similar quantities; see, for example, zhang2008improving, tsiatis2008covariate, jiang2022improving, jiang2022regression-adjusted and rafi2023efficient. We further illustrate that this feature of finely stratified experiments stems from the way in which finely stratified designs balance treatment status across covariate values, a property we define formally below and refer to as “fast-balancing.” Second, we derive a lower bound on the asymptotic variance of regular estimators of the parameter of interest in the form of a convolution theorem. This convolution theorem accommodates a large class of possible treatment assignment mechanisms, including covariate adaptive randomization (efron1971forcing; wei1978adaptive; zelen1974randomization; pocock1975sequential; hu2012asymptotic; bugni2018inference; ye2022inference; ma2020statistical; ma2024new), re-randomization (li2017general; li2018asymptotic; li2020rerandomization; li2020rerandomization-1; cytrynbaum2024finely), and fine stratification (jiang2021bootstrap; bai2022inference; cytrynbaum2023designing). We show that the lower bound is attained by the naïve method of moments estimator under a finely stratified design. In this sense, the naïve method of moments estimator under a finely stratified design is asymptotically efficient. More succinctly, we say that finely stratified experiments lead to efficient estimators “by design.” Finally, we strengthen this conclusion by characterizing all regular asymptotically linear estimators for a large class of treatment assignment mechanisms and use this characterization to establish conditions under which the fast-balancing property of finely stratified experiments is in fact a necessary condition for the naïve method of moments estimator to attain the efficiency bound.
@@
Together, these results demonstrate that finely stratified experiments lead to efficient estimators that prioritize transparency in that they preclude the researcher from “data snooping” associated with ex post nonparametric covariate adjustment. Importantly, concerns with this type of data snooping are not completely eliminated by typical pre-analysis plans because such adjustments involve choices, such as the choice of nonparametric estimator or tuning parameters, that are often not pre-registered prior to the experiment. The estimators are therefore attractive because they avoid performing nonparametric covariate adjustment in order to achieve efficiency and thereby remain “hands above the table” (freedman2008regression; lin2013agnostic).
@@
Our paper builds upon two strands of literature. The first strand of literature concerns the analysis of finely stratified experiments. Within this literature, our analysis is most closely related to bai2022inference, who derive the asymptotic behavior of the difference-in-means estimator of the ATE when treatment is assigned according to a matched pairs design, and cytrynbaum2023designing, who develops related results for an experimental design referred to as “local randomization” that permits the proportion of units assigned to treatment to vary with the baseline covariates. Beyond settings that study estimation of the ATE, bai2024inference-1 develops results for the analysis of different cluster-level average treatment effects and jiang2021bootstrap develop results analogous to those in bai2022inference for suitable estimators of the quantile treatment effect. This paper, like those just mentioned, operates in a “super-population” framework, in which the outcomes and covariates are assumed to be drawn as an i.i.d. sample from a population distribution. This is in contrast to an alternative strand of the literature that studies finely stratified experiments from the design-based perspective (see, in particular, imai2008variance; imai2009essential; fogarty2018mitigating; fogarty2018regression-assisted; liu2020regression-adjusted; pashley2021insights; bai2025new). To our knowledge, our paper is the first to analyze the properties of finely stratified experiments in a general framework that accommodates any parameter that can be characterized as the solution to a set of moment conditions involving a known function of the observed data. In more recent work, cytrynbaum2024finely considers a similar framework to study finely stratified re-randomized experiments, which nest finely stratified experiments as a special case. However, his results assume that the moment functions which define the parameter of interest are continuous in a way that precludes parameters like the Quantile Treatment Effect. Moreover, we emphasize that none of the above papers formally establish the asymptotic efficiency of finely stratified experiments. The second strand of literature concerns bounds on the efficiency with which treatment effect parameters can be estimated in experiments. We note that, due to the potential for dependence in treatment assignments across individuals, we cannot immediately appeal to standard semi-parametric efficiency results (see, for example, van_der_vaart1998asymptotic; chen2008semiparametric). Two important recent papers in this literature studying efficiency bounds in the special case of estimating the ATE are armstrong2022asymptotic and rafi2023efficient. Even in this special case, their results differ from ours in important and empirically relevant ways; Remark 4.3 provides an in-depth discussion of the connection between these results and ours. See also bai2022optimality for some finite-sample optimality properties of matched pairs designs for estimation of the ATE.
@@
The remainder of this paper is organized as follows. In Section 2, we describe our setup and notation. We emphasize, in particular, the way in which our framework can accommodate various treatment effect parameters of interest. Section 3 derives the asymptotic behavior of the naïve method of moments estimator of our parameter of interest when treatment is assigned using a finely stratified design and studies estimation of the asymptotic variance. In Section 4.1, we develop our lower bound on the asymptotic variance of regular estimators of these parameters and show that it is achieved by the the naïve method of moments estimator in a finely stratified design. In Section 4.2, we characterize all regular asymptotically linear estimators and argue that the fast-balancing property is necessary for the naïve method of moments estimator to attain the efficiency bound. In Section 5, we illustrate the practical relevance of our theoretical results through a simulation study. Finally, we conclude in Section 6 with some recommendations for empirical practice guided by both these simulations and our theoretical results. Proofs of all results can be found in the Appendix.
@@ 2 Setup and Motivation
We assume throughout that our sample consists of $n$ units. For any random vector indexed by $i$, for example $A_{i}$, we define $A^{(n)}=(A_{1},\ldots,A_{n})$. Let $P_{n}$ denote the distribution of the observed data $(R^{(n)},A^{(n)},X^{(n)})$, and $Q_{n}$ the distribution of $(R^{(n)}(1),R^{(n)}(0),X^{(n)})$. We assume $Q_{n}=Q^{n}$, where $Q$ is the marginal distribution of $(R_{i}(1),R_{i}(0),X_{i})$. Given $Q_{n}$, $P_{n}$ is then determined by (1) and the mechanism for determining treatment assignment. We assume that treatment assignment is performed such that a standard unconfoundedness assumptions holds and such that the probability of assignment given $X_{i}$ is some known constant for every $1\leq i\leq n$, as is often the case in most experiments:
@@
Assumption 2.1 restricts the probability of assignment to be the fixed fraction $\eta$ across the entire experimental sample, but this restriction can be weakened so that $\eta$ is replaced by $\eta(X_{i})$ for many of our subsequent results: see Remark 4.4 for a discussion. Given Assumption 2.1, it can be shown that $(X_{i},A_{i},R_{i})$ are identically distributed for $1\leq i\leq n$, and their marginal distribution does not change with $n$ (see Lemma A.7 in the Appendix). As a consequence, we denote the marginal distribution of $(X_{i},A_{i},R_{i})$ by $P$. We consider parameters $\theta_{0}\in\Theta\subset\mathbf{R}^{d_{\theta}}$ that can be defined as the solution to a set of moment equalities. As we show below, these parameters include a large class of causal parameters defined in terms of potential outcomes and potential treatments. Formally, let $m:\mathbb{R}^{d_{x}}\times\{0,1\}\times\mathbf{R}^{d_{r}}\to\mathbf{R}^{d_{\theta}}$ be a known measurable function, then we consider parameters $\theta_{0}$ that uniquely solve the moment equality
@@
For papers that consider estimators based on (4), see hirano2001estimation and hirano2003efficient.
@@
for $\theta=(\theta^{(1)},\theta^{(0)})^{\prime}$. Note that the quantile treatment effect $q_{Y(1)}(\tau)-q_{Y(0)}(\tau)$ can then be defined as $h(\theta_{0})$ where $h:\mathbb{R}^{2}\to\mathbb{R}$ is given by h(s,t)=sth(s,t)=s-t.
@@
Let $(\tilde{Y}_{i}(a),D_{i}(a))=R_{i}(a)$ denote the vector of potential outcomes ($\tilde{Y}_{i}(a)\in\mathbb{R}$) and treatment take-up ($D_{i}(a)\in\{0,1\}$) under treatment $a\in\{0,1\}$, and let $(Y_{i},D_{i})=R_{i}$ denote the vector of observed outcomes and treatment take-up. Note here that $\tilde{Y}_{i}(a)$ corresponds to the potential outcome under assignment $a\in\{0,1\}$ and not to the potential outcome for a given take-up $D_{i}=d$. Suppose $E_{Q}[D_{i}(1)-D_{i}(0)]\neq 0$ and let
@@
If we further assume instrument monotonicity (i.e., $P\{D_{i}(1)\geq D_{i}(0)\}=1$) and instrument exclusion, then $\theta_{0}$ could be re-interpreted as the local average treatment effect (LATE) in the sense of imbens1994identification.
@@
for some known function $\omega:\mathbb{R}^{d_{x}}\to\mathbb{R}$. It then follows from Assumption 2.1 that $\theta_{0}$ solves the moment condition in (3) with
@@
By defining $Y_{i}$ to be the average outcome in the $i$th cluster, $\theta_{0}$ defined in this way can accommodate the (cluster) size-weighted and equally-weighted average treatment effects considered in bugni2022inference and bai2024inference-1 in the context of cluster-level randomized controlled trials.
@@
where $\text{expit}(z)=\frac{\exp(z)}{1+\exp(z)}$. The log-odds ratio can then be defined as $h(\theta_{0})$ where $h:\mathbb{R}^{2}\to\mathbb{R}$ is given by $h(s,t)=t$. This parameter appears in, for example, zhang2008improving.
Additional examples could be obtained by considering combinations of Examples 2.12.5. For instance, combining the moment functions from Examples 2.3 and 2.4 would result in a weighted LATE parameter. Beyond these examples, certain treatment effect contrasts could also be related to the structural parameters in, for instance, an economic model of supply in demand: see, for example, the model estimated in casaburi2022using.
@@
Before proceeding, in the remainder of this section, we provide a more detailed summary of the main contributions of our paper. To this end, first note that if $A^{(n)}$ were assigned i.i.d., independently of $X^{(n)}$, then it can be shown under mild conditions on $m(\cdot)$ (see, for instance, Theorem 5.1 in van_der_vaart1998asymptotic) that the naïve method of moments estimator satisfies
@@
with $M=\frac{\partial}{\partial\theta^{\prime}}E_{P}[m(X,A,R,\theta)]\Big|_{\theta=\theta_{0}}$. In Section 3, we show that if we assign $A^{(n)}$ using a finely stratified design (see Assumption 3.1 below for a formal definition) then, under appropriate assumptions so that we achieve “fast balance” of the treatment across covariate values (see Assumptions 3.2, 3.3 below),
@@
where $\mu_{a}(X_{i})=E_{Q}[Y_{i}(a)|X_{i}]$. This choice of $m^{*}(\cdot)$ produces the well known doubly-robust moment condition for estimating the ATE (robins1995analysis; hahn1998role). It can then be shown that an appropriately constructed two-step estimator, in which $\mu_{1}(\cdot)$ and $\mu_{0}(\cdot)$ are non-parametrically estimated in a first step, attains $\mathbb{V}_{*}$ (tsiatis2008covariate; farrell2015robust; chernozhukov2017doubledebiasedneyman; rafi2023efficient). Intuitively, the estimator obtained from the augmented moment function $m^{*}(\cdot)$ performs nonparametric covariate adjustment by exploiting the information contained in $X^{(n)}$ that may not have been captured in the original moment function $m(\cdot)$. Similar nonparametric covariate adjustments based on augmented moment equations have been developed for other parameters of interest (zhang2008improving; belloni2017program; jiang2022improving; jiang2022regression-adjusted). In this sense, we show that fine stratification can perform nonparametric covariate adjustment “by design” for the large class of parameters that can be expressed in terms of moment conditions of the form given in (3), thus generalizing similar observations made in bai2022inference, bai2022optimality, and cytrynbaum2023designing in the special case of estimating the ATE. As we explain in the discussion following Theorem 3.1, this feature of finely stratified experiments is possible because fine stratification leads to fast-balancing of the treatment across covariate values, as defined formally in (14) below.
@@
Earlier work on efficient treatment effect estimation has noted that the variance $\mathbb{V}_{*}$ is in fact the efficiency bound for estimating $\theta_{0}$ under i.i.d. assignment (see, for instance, cattaneo2010efficient). A natural follow-up question is whether or not $\mathbb{V}_{*}$ continues to be the efficiency bound for estimating $\theta_{0}$ under a finely stratified design, or more generally for complex experimental designs which induce dependence in the treatment assignments across individuals in the experiment. In Section 4.1, we show that $\mathbb{V}_{\ast}$ continues to be the efficiency bound for estimating $\theta_{0}$ for a large class of treatment assignment mechanisms with a fixed marginal probability of treatment assignment, which includes finely stratified designs as a special case. We can thus conclude that, from the perspective of asymptotic efficiency, finely stratified designs are optimal experimental designs for a broad range of treatment effect estimation problems. In Section 4.2 we build on this result and establish conditions under which efficient estimation of $\theta_{0}$ using the naïve method of moments estimator can be achieved only if the experimental design leads to fast-balancing of the treatment across covariate values. In this sense, we show that the fast-balancing property of finely stratified experiments is in fact a necessary condition for achieving efficient estimation of $\theta_{0}$ “by design.”
3 The Asymptotic Variance of Finely Stratified Experiments

### v8

@@ Abstract
This paper studies the use of highly stratified designs for the efficient estimation of a large class of treatment effect parameters that arise in the analysis of experiments. By a “highly stratified” design, we mean experiments in which units are divided into blocks of a fixed size and a proportion within each block is assigned to a binary treatment uniformly at random. The class of parameters considered are those that can be expressed as the solution to a set of moment conditions constructed using a known function of the observed data. They include, among other things, average treatment effects, quantile treatment effects, and local average treatment effects as well as the counterparts to these quantities in experiments in which the unit is itself a cluster. In this setting, we establish three results. First, we show that under a highly stratified design, the naïve method of moments estimator achieves the same asymptotic variance as what could typically be attained under alternative treatment assignment mechanisms only through ex post covariate adjustment. Second, we argue that the naïve method of moments estimator under a highly stratified design is asymptotically efficient by deriving a lower bound on the asymptotic variance of regular estimators of the parameter of interest in the form of a convolution theorem. In this sense, highly stratified experiments are attractive because they lead to efficient estimators of treatment effect parameters “by design.” Finally, we strengthen this conclusion by establishing conditions under which a “fast-balancing” property of highly stratified designs is in fact necessary for the naïve method of moments estimator to attain the efficiency bound.
KEYWORDS: Convolution theorem, Efficiency, Experiment, Experimental design, Highly stratified experiment, Matched pairs, Randomized controlled trial
@@ 1 Introduction
This paper studies the use of highly stratified designs for the efficient estimation of a large class of treatment effect parameters that arise in the analysis of experiments. By a “highly stratified” design, we mean experiments in which units are divided into blocks of a fixed size based on their covariate values and a proportion within each group is assigned to a binary treatment uniformly at random. The canonical example of such a design is a matched pairs design: here the fixed size of the blocks equals two, and exactly one of the two units in each block is assigned to treatment at random, so that the marginal probability of treatment assignment is one half. More broadly, a “highly” stratified experiment is also intended to generalize the notion of a “finely” stratified experiment (as defined in Fogarty, 2018a), for which in every block there is exactly one treated or control individual. On the other hand, our use of the term “highly stratified” in this context contrasts with what we could call a “coarsely stratified” design, in which units are divided into a small set of large blocks: see Example 4.2 below for a formal definition. The class of parameters considered are those that can be expressed as the solution to a set of moment conditions constructed using a known function of the observed data. This class of parameters includes many causal parameters of interest: average treatment effects (ATEs), quantile treatment effects, and local average treatment effects as well as the counterparts to these quantities in experiments in which the unit is itself a cluster.
@@
In the setting described above, we establish three results. First, we study the asymptotic properties of a naïve method of moments estimator under a highly stratified design. Here, by a naïve method of moments estimator, we mean an estimator constructed using a direct sample analog of the moment conditions. For example, in the case of the ATE, such an estimator is given by the Horvitz-Thompson estimator for the difference in means. We show that under a highly stratified design, the naïve method of moments estimator achieves the same asymptotic variance as what could typically be attained under alternative treatment assignment mechanisms only through ex post covariate adjustment using the same set of covariates. Such adjustment strategies frequently involve the nonparametric estimation of conditional expectations or similar quantities; see, for example, Zhang et al. (2008), Tsiatis et al. (2008), Jiang et al. (2022a), Jiang et al. (2022b) and Rafi (2023). We further illustrate that this feature of highly stratified experiments stems from the way in which highly stratified designs balance treatment status across covariate values, a property we define formally below and refer to as “fast-balancing.” Second, we derive a lower bound on the asymptotic variance of regular estimators of the parameter of interest in the form of a convolution theorem. This convolution theorem accommodates a large class of possible treatment assignment mechanisms, including covariate adaptive randomization (Efron, 1971; Wei, 1978; Zelen, 1974; Pocock and Simon, 1975; Hu and Hu, 2012; Bugni et al., 2018; Ye et al., 2022; Ma et al., 2020, 2024), re-randomization (Li and Ding, 2017; Li et al., 2018; Li and Ding, 2020; Li et al., 2020; Cytrynbaum, 2024), and highly stratified designs (Jiang et al., 2021; Bai et al., 2022; Cytrynbaum, 2023b). We show that the lower bound is attained by the naïve method of moments estimator under a highly stratified design. In this sense, the naïve method of moments estimator under a highly stratified design is asymptotically efficient. More succinctly, we say that highly stratified experiments lead to efficient estimators “by design.” Finally, we strengthen this conclusion by characterizing all regular asymptotically linear estimators for a large class of treatment assignment mechanisms and use this characterization to establish conditions under which the fast-balancing property of highly stratified experiments is in fact a necessary condition for the naïve method of moments estimator to attain the efficiency bound.
@@
Together, these results demonstrate that highly stratified experiments lead to efficient estimators that prioritize transparency in that they preclude the researcher from “data snooping” associated with ex post nonparametric covariate adjustment. Importantly, concerns with this type of data snooping are not completely eliminated by typical pre-analysis plans because such adjustments involve choices, such as the choice of nonparametric estimator or tuning parameters, that are often not pre-registered prior to the experiment. The estimators are therefore attractive because they avoid performing nonparametric covariate adjustment in order to achieve efficiency and thereby remain “hands above the table” (Freedman, 2008; Lin, 2013).
@@
Our paper builds upon two strands of literature. The first strand of literature concerns the analysis of highly stratified experiments. Within this literature, our analysis is most closely related to Bai et al. (2022), who derive the asymptotic behavior of the difference-in-means estimator of the ATE when treatment is assigned according to a matched pairs design, and Cytrynbaum (2023b), who develops related results for an experimental design referred to as “local randomization” that permits the proportion of units assigned to treatment to vary with the baseline covariates. Beyond settings that study estimation of the ATE, Bai et al. (2024b) develops results for the analysis of different cluster-level average treatment effects and Jiang et al. (2021) develop results analogous to those in Bai et al. (2022) for suitable estimators of the quantile treatment effect. This paper, like those just mentioned, operates in a “super-population” framework, in which the outcomes and covariates are assumed to be drawn as an i.i.d. sample from a population distribution. This is in contrast to an alternative strand of the literature that studies highly stratified experiments from the design-based perspective (see, in particular, Imai, 2008; Imai et al., 2009; Fogarty, 2018a, b; Liu and Yang, 2020; Pashley and Miratrix, 2021; Bai et al., 2025b). To our knowledge, our paper is the first to analyze the properties of highly stratified experiments in a general framework that accommodates any parameter that can be characterized as the solution to a set of moment conditions involving a known function of the observed data. In more recent work, Cytrynbaum (2024) considers a similar framework to study highly stratified re-randomized experiments, which nest highly stratified experiments as a special case. However, his results assume that the moment functions which define the parameter of interest are continuous in a way that precludes parameters like the Quantile Treatment Effect. Moreover, we emphasize that none of the above papers formally establish the asymptotic efficiency of highly stratified experiments. The second strand of literature concerns bounds on the efficiency with which treatment effect parameters can be estimated in experiments. We note that, due to the potential for dependence in treatment assignments across individuals, we cannot immediately appeal to standard semi-parametric efficiency results (see, for example, van der Vaart, 1998; Chen et al., 2008). Two important recent papers in this literature studying efficiency bounds in the special case of estimating the ATE are Armstrong (2022) and Rafi (2023). Even in this special case, their results differ from ours in important and empirically relevant ways; Remark 4.3 provides an in-depth discussion of the connection between these results and ours. See also Bai (2022) for some finite-sample optimality properties of matched pairs designs for estimation of the ATE.
@@
The remainder of this paper is organized as follows. In Section 2, we describe our setup and notation. We emphasize, in particular, the way in which our framework can accommodate various treatment effect parameters of interest. Section 3 derives the asymptotic behavior of the naïve method of moments estimator of our parameter of interest when treatment is assigned using a highly stratified design and studies estimation of the asymptotic variance. In Section 4.1, we develop our lower bound on the asymptotic variance of regular estimators of these parameters and show that it is achieved by the the naïve method of moments estimator in a highly stratified design. In Section 4.2, we characterize all regular asymptotically linear estimators and argue that the fast-balancing property is necessary for the naïve method of moments estimator to attain the efficiency bound. In Section 5, we illustrate the practical relevance of our theoretical results through a simulation study. Finally, we conclude in Section 6 with some recommendations for empirical practice guided by both these simulations and our theoretical results. Proofs of all results can be found in the supplementary material.
@@ 2 Setup and Motivation
We assume throughout that our sample consists of $n$ units. For any random vector indexed by $i$, for example $A_{i}$, we define $A^{(n)}=(A_{1},\ldots,A_{n})$. Let $P_{n}$ denote the distribution of the observed data $(R^{(n)},A^{(n)},X^{(n)})$, and $Q_{n}$ the distribution of $(R^{(n)}(1),R^{(n)}(0),X^{(n)})$. We assume $Q_{n}=Q^{n}$, where $Q$ is the marginal distribution of $(R_{i}(1),R_{i}(0),X_{i})$. Given $Q_{n}$, $P_{n}$ is then determined by (1) and the mechanism for determining treatment assignment. We assume that treatment assignment is performed such that a standard unconfoundedness assumption holds and such that the probability of assignment given $X_{i}$ is some known constant for every $1\leq i\leq n$, as is often the case in most experiments:
@@
Assumption 2.1 restricts the probability of assignment to be the fixed fraction $\eta$ across the entire experimental sample, but this restriction can be weakened so that $\eta$ is replaced by $\eta(X_{i})$ for many of our subsequent results: see Remark 4.4 for a discussion. Given Assumption 2.1, it can be shown that $(X_{i},A_{i},R_{i})$ are identically distributed for $1\leq i\leq n$, and their marginal distribution does not change with $n$ (see Lemma A.7 in the Supplement). As a consequence, we denote the marginal distribution of $(X_{i},A_{i},R_{i})$ by $P$. We consider parameters $\theta_{0}\in\Theta\subset\mathbf{R}^{d_{\theta}}$ that can be defined as the solution to a set of moment equalities. As we show below, these parameters include a large class of causal parameters defined in terms of potential outcomes and potential treatments. Formally, let $m:\mathbf{R}^{d_{x}}\times\{0,1\}\times\mathbf{R}^{d_{r}}\to\mathbf{R}^{d_{\theta}}$ be a known measurable function, then we consider parameters $\theta_{0}$ that uniquely solve the moment equality
@@
For papers that consider estimators based on (4), see Hirano and Imbens (2001) and Hirano et al. (2003).
@@
for $\theta=(\theta^{(1)},\theta^{(0)})^{\prime}$. Note that the quantile treatment effect $q_{Y(1)}(\tau)-q_{Y(0)}(\tau)$ can then be defined as $h(\theta_{0})$ where $h:\mathbf{R}^{2}\to\mathbf{R}$ is given by h(s,t)=sth(s,t)=s-t.
@@
Let $(\tilde{Y}_{i}(a),D_{i}(a))=R_{i}(a)$ denote the vector of potential outcomes ($\tilde{Y}_{i}(a)\in\mathbf{R}$) and treatment take-up ($D_{i}(a)\in\{0,1\}$) under treatment $a\in\{0,1\}$, and let $(Y_{i},D_{i})=R_{i}$ denote the vector of observed outcomes and treatment take-up. Note here that $\tilde{Y}_{i}(a)$ corresponds to the potential outcome under assignment $a\in\{0,1\}$ and not to the potential outcome for a given take-up $D_{i}=d$. Suppose $E_{Q}[D_{i}(1)-D_{i}(0)]\neq 0$ and let
@@
If we further assume instrument monotonicity (i.e., $P\{D_{i}(1)\geq D_{i}(0)\}=1$) and instrument exclusion, then $\theta_{0}$ could be re-interpreted as the local average treatment effect (LATE) in the sense of Imbens and Angrist (1994).
@@
for some known function $\omega:\mathbf{R}^{d_{x}}\to\mathbf{R}$. It then follows from Assumption 2.1 that $\theta_{0}$ solves the moment condition in (3) with
@@
By defining $Y_{i}$ to be the average outcome in the $i$th cluster, $\theta_{0}$ defined in this way can accommodate the (cluster) size-weighted and equally-weighted average treatment effects considered in Bugni et al. (2022) and Bai et al. (2024b) in the context of cluster-level randomized controlled trials.
@@
where $\text{expit}(z)=\frac{\exp(z)}{1+\exp(z)}$. The log-odds ratio can then be defined as $h(\theta_{0})$ where $h:\mathbf{R}^{2}\to\mathbf{R}$ is given by $h(s,t)=t$. This parameter appears in, for example, Zhang et al. (2008).
Additional examples could be obtained by considering combinations of Examples 2.12.5. For instance, combining the moment functions from Examples 2.3 and 2.4 would result in a weighted LATE parameter. Beyond these examples, certain treatment effect contrasts could also be related to the structural parameters in, for instance, an economic model of supply and demand: see, for example, the model estimated in Casaburi and Reed (2022).
@@
Before proceeding, in the remainder of this section, we provide a more detailed summary of the main contributions of our paper. To this end, first note that if $A^{(n)}$ were assigned i.i.d., independently of $X^{(n)}$, then it can be shown under mild conditions on $m(\cdot)$ (see, for instance, Theorem 5.1 in van der Vaart, 1998) that the naïve method of moments estimator satisfies
@@
with $M=\frac{\partial}{\partial\theta^{\prime}}E_{P}[m(X,A,R,\theta)]\Big|_{\theta=\theta_{0}}$. In Section 3, we show that if we assign $A^{(n)}$ using a highly stratified design (see Assumption 3.1 below for a formal definition) then, under appropriate assumptions so that we achieve “fast balance” of the treatment across covariate values (see Assumptions 3.2, 3.3 below),
@@
where $\mu_{a}(X_{i})=E_{Q}[Y_{i}(a)|X_{i}]$. This choice of $m^{*}(\cdot)$ produces the well known doubly-robust moment condition for estimating the ATE (Robins et al., 1995; Hahn, 1998). It can then be shown that an appropriately constructed two-step estimator, in which $\mu_{1}(\cdot)$ and $\mu_{0}(\cdot)$ are non-parametrically estimated in a first step, attains $\mathbb{V}_{*}$ (Tsiatis et al., 2008; Farrell, 2015; Chernozhukov et al., 2017; Rafi, 2023). Intuitively, the estimator obtained from the augmented moment function $m^{*}(\cdot)$ performs nonparametric covariate adjustment by exploiting the information contained in $X^{(n)}$ that may not have been captured in the original moment function $m(\cdot)$. Similar nonparametric covariate adjustments based on augmented moment equations have been developed for other parameters of interest (Zhang et al., 2008; Belloni et al., 2017; Jiang et al., 2022a, b). In this sense, we show that highly stratified designs can perform nonparametric covariate adjustment “by design” for the large class of parameters that can be expressed in terms of moment conditions of the form given in (3), thus generalizing similar observations made in Bai et al. (2022), Bai (2022), and Cytrynbaum (2023b) in the special case of estimating the ATE. As we explain in the discussion following Theorem 3.1, highly stratified experiments have this feature because they lead to fast-balancing of the treatment across covariate values, as defined formally in (14) below.
@@
Earlier work on efficient treatment effect estimation has noted that the variance $\mathbb{V}_{*}$ is in fact the efficiency bound for estimating $\theta_{0}$ under i.i.d. assignment (see, for instance, Cattaneo, 2010). A natural follow-up question is whether or not $\mathbb{V}_{*}$ continues to be the efficiency bound for estimating $\theta_{0}$ under a highly stratified design, or more generally for complex experimental designs which induce dependence in the treatment assignments across individuals in the experiment. In Section 4.1, we show that $\mathbb{V}_{\ast}$ continues to be the efficiency bound for estimating $\theta_{0}$ for a large class of treatment assignment mechanisms with a fixed marginal probability of treatment assignment, which includes highly stratified designs as a special case. We can thus conclude that, from the perspective of asymptotic efficiency, highly stratified designs are optimal experimental designs for a broad range of treatment effect estimation problems. In Section 4.2 we build on this result and establish conditions under which efficient estimation of $\theta_{0}$ using the naïve method of moments estimator can be achieved only if the experimental design leads to fast-balancing of the treatment across covariate values. In this sense, we show that the fast-balancing property of highly stratified experiments is in fact a necessary condition for achieving efficient estimation of $\theta_{0}$ “by design.”
3 The Asymptotic Variance of Highly Stratified Experiments